A block of a material has a Volume of 20cm cube and a density of 2.5g/cm cube. It is suspended from a spring balance with half of the volume of the block immersed in water . What is the reading of the spring balance?
1
Expert's answer
2020-08-27T10:24:52-0400

"m=\frac{F}{g}=m-0.5\rho V=20\cdot2.5-0.5\cdot1\cdot20=50-10=40(g)"